period P = \frac{1}{f}

t_{1} = \frac{L}{v}

t_{2} = \frac{L}{v}

t_{roundtrip} = t_{1} + t_{2} = \frac{2L}{v}

y_{1}(x, t) = A sin(kx + ωt)

y_{2}(x, t) = A sin(kx − ωt)

sin(α) + sin(β) = 2 sin (\frac{α + β}{2} ) cos (\frac{α − β}{2} )

v = \sqrt{\frac{T}{μ}}